# Experimental Study of Bremsstrahlung Gamma Ray Emission and Short-Range Correlations in 124Sn+124Sn Collisions at 25 MeV/u

### v1

@@ V.3 Experimental $\gamma$ spectrum in fast coincidence
The same comparison between the background-subtracted $\gamma$ spectrum (excluding the SSD M2  trigger) and theoretical model predictions, scaled by a factor of $1.7\times 10^{9}$, is shown in Fig. 28. Using a likelihood-based method, we compare the experimental spectrum with IBUU model predictions and extract an optimal $R_{\text{HMT}}=(18.0\pm 2.8)\%$. This value is consistent, within uncertainties, with the result obtained via beam-off background subtraction method, indicating that the extracted SRC fraction is robust and independent of the specific background subtraction method.
@@
VI Reconstruction of the Original $\gamma$ Spectrum
Finally, in order to offer a $\gamma$ spectrum which can be compared to other theoretic calculations without knowing the response matrix, we also tried to reconstructed the original $\gamma$ spectrum from the measured one by solving an inverse problem. In addition, the results of $R_{\rm HMT}$ can also be checked one more time from the original spectrum. For this purpose, the Richardson Lucy (RL) algorithm, which has been widely applied in the nuclear physics [92, 93, 94, 95, 96, 97, 98], is adopted to solve the inverse problem [65].
VI.1 Richardson Lucy Algorithm
@@
where $P(\nu|\mu)$ represents the conditional probability of measuring a photon with real property $\mu$ as $\nu$.
@@
To assess statistical uncertainties, we perform multinomial sampling of the measured spectrum. Multiple pseudo-spectra are generated while preserving the total event count, and each is independently processed through the RL algorithm. The resulting ensemble of reconstructed spectra allows us to compute the standard deviation for each energy bin, which serves as our uncertainty estimate. In the RL reconstructed $\gamma$-ray spectrum, the central values are obtained by the measured spectrum, while the error bars reflect the propagated statistical uncertainties derived from the multinomial sampling. This method robustly accounts for statistical fluctuations in the measurement and their impact on the spectral reconstruction.
VI.2 Original Bremsstrahlung $\gamma$ Spectrum
The initial values for all iterative solutions of different spectra are set to the theoretical curve with $R_{\text{HMT}}=0\%$. This choice eliminates any prior bias toward a non-zero HMT ratio in the initial condition. By starting with $R_{\text{HMT}}=0\%$, the iterative process avoids the risk of a pre-imposed HMT contribution influencing the reconstructed spectrum. Furthermore, this initialization improves both the efficiency and accuracy of the iterations, ensuring that the final solution is derived purely from the input data without relying on an arbitrary starting condition.
@@ VI.2 Original Bremsstrahlung $\gamma$ Spectrum
where $e_{i}$ represents the input measured value for the $i$th energy bin, $e_{i}^{P}$ is the predicted value from the current iteration, and $N_{p}$ is the total number of bins. Here, $N_{p}=100$, corresponding to the energy range of $1~\text{MeV}$ to $100~\text{MeV}$ with a step size of $1~\text{MeV}$.
@@
The evolution of $\delta_{\text{exp}}$ as a function of iteration number in solving central spectrum is shown in Fig. 29(a). It can be observed that $\delta_{\text{exp}}$ gradually decreases with increasing iteration times, indicating a systematic improvement in the agreement between the experimental and predicted spectra. The end of curve is at the convergence point, signaling the optimal termination of the iterative process.
@@
After the iteration process described earlier, the reconstructed real $\gamma$ spectrum from the central measured spectrum is obtained and shown as black hollow dots in Fig. 30. The spectrum has been transformed from the lab frame to the center-of-mass frame and subsequently rebinned, combining every 4 bins into 1.
@@
To present the comparison more clearly, the main theoretical curves are scaled by a specific coefficient of $4.5\times 10^{9}$ and rebin number to match the experimental data format. The comparison between the reconstructed spectrum and the scaled theoretical curves is shown in Fig. 30.
@@
VI.3 Direct Comparison to IBUU-MDI simulations
@@
$\chi^{2}(R_{\text{HMT}})=\sum_{i}\frac{1}{\sigma_{i}^{2}}(\text{counts}_{i}^{\text{exp}}-\text{counts}_{i}^{R_{\text{HMT}}})^{2},$ (35)
where $\text{counts}_{i}^{\text{exp}}$ and $\text{counts}_{i}^{R_{\text{HMT}}}$ represent the counts in the $i$th energy bin for the experimental data and the theoretical prediction with a given $R_{\text{HMT}}$, respectively, and $\sigma_{i}$ is the uncertainty of the reconstructed experimental spectrum in the $i^{\rm th}$ bin.
In addition, due to the need to compare the shape of the experimental data with the theoretical predictions, the theoretical curves must be scaled again when calculating the $\chi^{2}(R_{\text{HMT}})$. This scaling ensures that we minimize the value of $\chi^{2}(R_{\text{HMT}})$ for each theoretical curve, providing an accurate evaluation of the best match between the experimental and theoretical spectra. This procedure allows us to determine the optimal $R_{\text{HMT}}$ that best describes the data while accounting for differences in the overall amplitude between the experimental and theoretical curves. The theoretical curve that minimizes $\chi^{2}(R_{\text{HMT}})$ corresponds to the best-fit $R_{\text{HMT}}$, providing the most accurate description of the experimental data.
@@
VII Conclusion and Outlook
In conclusion, SRC in atomic nuclei has drawn increasing attention in nuclear physics, since it is not only relavant to the nuclear structure and the fundamental implication of the nuclear force, but also essential to understand the dense nuclear matter existing in nuclear stars. Moerover, SRC is also related to modification of the parton distribution of nucleons in atomic nuclei (nPDF) according to the QCD factorization scheme. While enormous progress has been made in electron-nucleus scattering experiment and nucleus-proton knock-out reactions, experimental progress in HICs is still scarce.
The bremsstrahlung $\gamma$ emissions in HICs has been identified as a novel probe to detect the SRC because of the fact that the HMT of the nucleons arising from SRC hardens the high energy spectrum of the bremsstrahlung $\gamma$-rays. Following this suggestions, two experiments has been performed on CSHINE. The previous experiment of 86Kr+124Sn at 25 MeV/u suggested the existence of HMT, offering the motivation to detect the SRC using bremsstrahlung $\gamma$ emissions.
@@ VII Conclusion and Outlook
To be more quantitative and more precise, the new experiment of 124Sn+124Sn at 25 MeV/u has been performed with much improvements on the experimental setup as well as data analysis, including the installation of the veto scintillators surrounding the CsI(Tl) hodoscope and the extension of the energy range compared to the previous one, as well as the test of the non-linearity of the response of the CsT(Tl) crystal using the quasi monochromatic $\gamma$ beam at SLEGS. The total statistics has been improved. Moreover, the adoption of a symmetric system in the experiment allows to determine the $R_{\rm HMT}$ to the specific nucleus 124Sn.
@@
Two approaches have been applied to obtain the bremsstrahlung $\gamma$-ray energy spectrum. The main approach is to use the whole data set in a slow coincidence gate by subtracting the cosmic-ray muon background which has been measured with beam-off. The other one as a cross check is to use the data in a fast coincidence gate, while the background is obtained in a beam-on random coincidence time window with the same width. With the second approach, the SSD M2  trigger condition is necessarily excluded since the trigger signal suffers seriously time jitter. Both approaches reach the consistent constraint on the HMT fraction within uncertainty. Combined, the fraction of high-momentum nucleons arising from SRC has been derived as $R_{\rm HMT}=(20\pm 3)\%$.
@@
From the experimentally measured spectrum, we also reconstructed the original bremsstrahlung $\gamma$-ray energy spectrum by applying the Richardson-Lucy deblurring algorithm, which is originated from the application of optic deblurring method to solve the inverse problem. The inference of the fraction of the HMT nucleons is in accordance with the above result. Advantageously, the reconstruction of the original $\gamma$ spectrum allows for direct comparison to theoretic calculations without knowing the detector response matrix by other theoretic groups in the future.
@@
The current experimental result gives a precise fraction of SRC nucleons in nuclei, confirming the nonzero $R_{\rm HMT}$ beyond $5\sigma$ confidential level. For the first time, our work establishes a new paradigm to probe the SRC and the parton dynamics in nuclei using low-energy heavy-ion collisions.
@@
Acknowledgement  This work is supported by the National Natural Science Foundation of China under Grant Nos. 12335008, 12205160, 12275129, by the Ministry of Science and Technology under Grant No. 2022YFE0103400, by the Center for High Performance Computing and Initiative Scientific Research Program in Tsinghua University and by HIRFL. The authors acknowledge the HIRFL machine team for providing the excellent beam condition.

### v2

@@ IV.3 Experimental $\gamma$ spectrum in fast coincidence
The same comparison between the background-subtracted $\gamma$ spectrum (excluding the SSD M2  trigger) and theoretical model predictions, scaled by a factor of $1.7\times 10^{9}$, is shown in Fig. 8. Using a likelihood-based method, we compare the experimental spectrum with IBUU model predictions and extract an optimal $R_{\text{HMT}}=(18.0\pm 2.8)\%$. This value is consistent, within uncertainties, with the result obtained via the beam-off background subtraction method, indicating that the extracted SRC fraction is robust and independent of the specific background subtraction method.
@@
V Reconstruction of the Original $\gamma$ Spectrum
Finally, in order to offer a $\gamma$ spectrum that can be compared with other theoretic calculations without knowing the response matrix, we also tried to reconstruct the original $\gamma$ spectrum from the measured one by solving an inverse problem. In addition, the results of $R_{\rm HMT}$ can also be checked one more time from the original spectrum. For this purpose, the Richardson Lucy (RL) algorithm, which has been widely applied in nuclear physics [79, 14, 60, 58, 52, 96, 59], is adopted to solve the inverse problem [95].
V.1 Unfolding Method and Detector Response
@@
where $P(\nu|\mu)$ represents the conditional probability of measuring a photon with the real property $\mu$ as $\nu$.
@@
To assess statistical uncertainties, we performed multinomial sampling of the measured spectrum. Multiple pseudo-spectra are generated while preserving the total event count, and each is independently processed through the RL algorithm. The resulting ensemble of reconstructed spectra allows us to compute the standard deviation for each energy bin, which serves as our uncertainty estimate. In the RL reconstructed $\gamma$-ray spectrum, the central values are obtained by the measured spectrum, while the error bars reflect the propagated statistical uncertainties derived from the multinomial sampling. This method robustly accounts for statistical fluctuations in the measurement and their impact on the spectral reconstruction.
V.2 Original $\gamma$-Ray Spectrum
The initial values for all iterative solutions of different spectra are set to the theoretical curve with $R_{\text{HMT}}=0\%$. This choice eliminates any prior bias toward a nonzero HMT ratio in the initial condition. By starting with $R_{\text{HMT}}=0\%$, the iterative process avoids the risk of a pre-imposed HMT contribution influencing the reconstructed spectrum. Furthermore, this initialization improves both the efficiency and accuracy of the iterations, ensuring that the final solution is derived purely from the input data without relying on an arbitrary starting condition.
@@ V.2 Original $\gamma$-Ray Spectrum
where $e_{i}$ represents the input measured value for the $i^{\rm th}$ energy bin, $e_{i}^{P}$ is the predicted value from the current iteration, and $N_{p}$ is the total number of bins. Here, $N_{p}=100$, corresponding to the energy range of $1~\text{MeV}$ to $100~\text{MeV}$ with a step size of $1~\text{MeV}$.
@@
The evolution of $\delta_{\text{exp}}$ as a function of iteration number in solving the central spectrum is shown in Fig. 9(a). It can be observed that $\delta_{\text{exp}}$ gradually decreases with increasing iteration times, indicating a systematic improvement in the agreement between the experimental and predicted spectra. The end of the curve is at the convergence point, signaling the optimal termination of the iterative process.
@@
After the iteration process described earlier, the reconstructed real $\gamma$ spectrum from the central measured spectrum is obtained and shown as black dots in Fig. 10. The spectrum has been transformed from the lab frame to the center-of-mass frame and subsequently rebinned, combining every 4 bins into 1.
@@
To present the comparison more clearly, the main theoretical curves are scaled by a specific coefficient of $4.5\times 10^{9}$ and a rebin number to match the experimental data format. The comparison between the reconstructed spectrum and the scaled theoretical curves is shown in Fig. 10.
@@
V.3 Direct Comparison with IBUU-MDI Calculations
@@
$\chi^{2}(R_{\text{HMT}})=\sum_{i}\frac{1}{\sigma_{i}^{2}}(\text{S}_{i}^{\text{exp}}-\alpha\text{S}_{i}^{R_{\text{HMT}}})^{2},$ (24)
where $\text{S}_{i}^{\text{exp}}$ and $\text{S}_{i}^{R_{\text{HMT}}}$ represent the counts in the $i^{\rm th}$ energy bin for the experimental data and the theoretical prediction with a given $R_{\text{HMT}}$, respectively, and $\sigma_{i}$ is the uncertainty of the reconstructed experimental spectrum in the $i^{\rm th}$ bin. The parameter $\alpha$ is a normalization factor determined by minimizing $\chi^{2}$ for each theoretical curve.
Since the comparison focuses on the spectral shape rather than the absolute yield, the normalization factor only adjusts the overall amplitude of the theoretical spectrum. Therefore, this scaling does not influence the extracted value of $R_{\text{HMT}}$, which is determined by the differences in spectral shape. The theoretical curve that minimizes $\chi^{2}(R_{\text{HMT}})$ corresponds to the best-fit $R_{\text{HMT}}$, providing the most accurate description of the experimental data.
@@
VI Conclusion and Outlook
@@
In conclusion, the Bremsstrahlung $\gamma$-ray emissions were measured with high precision in 124Sn+124Sn at 25 MeV/u using the experiment CSHINE. In comparison to the previous experiment 86Kr+124Sn, significant advancements were made in both the experimental setup and data analysis. These improvements include the installation of veto scintillators surrounding the CsI(Tl) hodoscope, the extension of the energy range compared to earlier experiments, and the testing of the non-linearity of the CsI(Tl) crystal response using quasi-monochromatic $\gamma$-rays from the Shanghai Laser Electron Gamma Source. The total statistics were also enhanced. Furthermore, the choice of symmetric reaction system enabled us to determine the HMT fraction ($R_{\rm HMT}$) for the specific nucleus 124Sn.
@@
Two methods were employed to obtain the bremsstrahlung $\gamma$-ray energy spectrum. The primary method involved using the entire data set within a slow coincidence gate, with the cosmic-ray muon background subtracted, as measured with the beam off. The secondary method, used as a closure check, utilized data from a fast coincidence gate, with the background obtained from a beam-on random coincidence time window of the same width. The second method necessarily excluded the SSD M2  trigger signal, as it is suffered from significant time jitter. Both approaches provided consistent constraints on the HMT fraction within the measurement uncertainties. From these analyses, we derived the fraction of high-momentum nucleons arising from SRC to be $R_{\rm HMT}=(20\pm 3)\%$.
@@
From the experimentally measured spectrum, we also reconstructed the original bremsstrahlung $\gamma$-ray energy spectrum by applying the Richardson-Lucy deblurring algorithm, which originates from the application of the optical deblurring method to solve the inverse problem. The fraction of HMT nucleons inferred from this reconstruction was consistent with the result obtained above. This reconstruction of the original $\gamma$-ray spectrum enables direct comparisons with future theoretical calculations, independent of the detector response matrix.
@@
The current experimental result provides a statistically precise determination of the SRC nucleon fraction in nuclei, confirming a non-zero $R_{\rm HMT}$ at a significance level greater than $5\sigma$. For the first time, our work establishes a new and complementary approach to probe SRC and parton dynamics in nuclei via low-energy heavy-ion collisions.
@@
Acknowledgement  This work is supported by the National Natural Science Foundation of China under Grant Nos. 12335008, 12205160, 12275129, by the Ministry of Science and Technology under Grant No. 2022YFE0103400, by the Center for High Performance Computing and the Initiative Scientific Research Program in Tsinghua University and by HIRFL. The authors acknowledge the HIRFL machine team for providing excellent beam conditions.